# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ Abstract
Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades. Our new near-optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound. Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients. Due to this feature our root-finders (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for a polynomial that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence. Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large. Our auxiliary algorithms and techniques for computations with black box polynomials can be of independent interest.
@@
PART I: BACKGROUND AND OUR PROGRESS
@@ 1 State of the Art and our Main Results
In this part of the paper we state our main results and outline our algorithms. We present them in some detail in Parts II, III, and IV, which can be read independently of each other.
@@ 1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2$, according to Galois theory of 1830, 1832.
Nevertheless, in 1924 in [149], Herman Weyl published his constructive proof of the Fundamental Theorem of Algebra where he rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough. This was the solution of the following Problem 0 of modern polynomial root-finding:
@@
a positive error tolerance $\epsilon=1/2^{b}>0$, an integer $m$, $1\leq m\leq d$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle, such that $z_{j}\in\mathbb{D}$ if and only if $j\leq m$.
OUTPUT: approximations $\tilde{z}_{j}\approx z_{j}$, $j=1,\dots,m$, to the $m$ zeros of $p(x)$ (aka roots of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$. lying in the domain $\mathbb{D}$ such that
$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm for}~{}j=1,\dots,m.$ (1.2)
@@
[Approximation of all roots.] Fix $R\geq\max_{j}|z_{j}|$ (see [52] for simple expressions for such $R$) and let $\mathbb{D}$ denote the disc $D(0,R):=\{x:~{}|x\leq R|\}$, centered in the origin and having radius $R$. Then Problem 0 requires us to approximate all $d$ roots.
@@
[Numerical approximation of real roots.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots (cf. Kac [66], Edelman and Kostlan [39]). More generally consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon/\sqrt{2}$-neighborhood of a fixed real line segment. To solve this problem we should approximate both real and near-real roots, that is, their perturbations due to small input and rounding errors. The known real root-finders (cf. [80, Chs. 2 and 3]) do not address this important issue.
@@
Major motivation for modern study of Problem 0 comes from application to approximation of the eigenvalues of a matrix (see our Secs. 1.3 and 1.5) and the solution of a multivariate polynomial system of equations [29], both problems being highly important for the theory and practice of modern computations. [82, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics. Furthermore, Problem 0 for a domain $\mathbb{D}$ containing all roots is closely linked to polynomial factorization (see Secs. 24 and 25), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147].
@@
Hundreds of efficient root-finders have been proposed [80, 82] and keep appearing (see [122, 22, 40, 70, 76, 127, 108, 23, 24, 64, 110, 111, 75, 60, 62, 81, 63, 59, 115], and the bibliography therein). In [133] Smale proposed to study Problem 0 in terms of arithmetic complexity and computational precision and later, jointly with Blum, Cucker, and Shub (cf. [14, 30]), linked this study to rounding errors and conditioning of the roots. In [134] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of the number of bit operations involved and presented an algorithm that solved Problem 0 for a domain that contains all roots by using $\tilde{O}(bd^{3})$ bit operations, [^2] Here and hereafter we assume that $b\geq d$, recalling that error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable. that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$.
@@
His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [96] (also see its journal versions [97, 103] and surveys [99, 100]); it supports bound $\tilde{O}(bd^{2})$ and showed the information lower bound $0.25(d+1)bd$ (see Cor. 24.2 in Sec. 24.2).
If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated, as we specify in Def. 24.1 of Sec. 24, then the lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$ (see Sec. 24), and so do the upper bounds supported by the algorithm of [96, 97, 103] as well; hence the algorithm remains near-optimal (see Thms. 24.1 and 24.4 and Cor. 24.3 in Sec. 24).
@@ 1.2 Reduction to the case of the unit disc
Hereafter T shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively. Furthermore, let Problem $\widehat{0}$ denote Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
@@
(i) Two ways from the roots in the unit disc to all roots. Given a value
$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.4)
(see [52] for simple expressions for such $R$), we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
Schönhage in [134] reduced approximation of all $d$ roots to Problem $\widehat{0}$ without assuming (1.4) or using (1.3) as follows. Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
@@
(ii) From roots in the unit disc to roots in a bounded domain. Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.3), and finally, (d) use the map (1.3) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$. In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$; in Sec. 2.2 we control the computational cost of that additional work.
1.3 Black box root-finders; the isolation of roots, eigenvalues, and their clusters
A nontrivial variant of Newton’s iterations by Louis and Vempala [76], at FOCS 2016, approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.4). While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [76] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation. [76] shows benefits of extension of such algorithms to approximation of the eigenvalues of a matrix. Next we state the relevant tasks.
@@ 1.3 Black box root-finders; the isolation of roots, eigenvalues, and their clusters
Problem 2: Approximation of the Eigenvalues of a Matrix. [^3] As in [76], the problem covers or can be extended to the case where an input is a polynomial matrix, aka matrix polynomial.
INPUT: $\epsilon=1/2^{b}>0$, integers $m$ and $d$ such that $0<m\leq d$, and a $d\times d$ matrix $M$, and a complex domain $\mathbb{D}$ that contains precisely $m$ eigenvalues of $M$ counted with their multiplicities.
OUTPUT: Approximations within $\epsilon$ to each of all $m$ eigenvalues of $M$ lying in the domain $\mathbb{D}$.
@@
Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of a polynomial and the Eigenvalues of a matrix. Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all roots or eigenvalues that lie in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
@@ 1.4.1 Why we should use them
One can perform black box evaluation at $q>d$ points $x_{1},\dots,x_{q}$ and then interpolate to $p$. The bit operation cost of interpolation, however, is unbounded unless the $q$ points are well isolated from the roots, but it is a challenge to yield such isolation. Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1 ,1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots.
Moreover, here are three examples of important benefits of using black box algorithms even where an input polynomial is given with its coefficients.
(i) Application of a black box root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ matrix $M$ only involves the values of this polynomial but not its coefficients, whose overall size exceeds the size (a norm) of the matrix $M$ by a factor of $d$. Consequently, extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [134, Sec. 21].
@@
(iii) Black box polynomial root-finders are particularly efficient where a $d$th degree polynomial $p$ can be evaluated fast, say, in $O(\log(d))$ arithmetic operations (hereafter (ops) such as a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$. For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [89, 69] or [139, Section “Pan’s method"]).
1.4.2 What we adopt and what cannot adopt for black box polynomial root-finding

### v37

# Polynomial Root-Finding and Algebraic Eigenvalue Problem

@@ Abstract
Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades. Our new near-optimal root-finders approximate all complex zeros of a polynomial almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound. The root-finders work even for a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients. Hence they (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for polynomials that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence. Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large. We elaborate upon design and analysis of our algorithms, comment on their potential heuristic acceleration, and briefly cover polynomial root-finding by means of functional iterations. Our techniques can be of independent interest.
@@
PART I: State of the Art and Our Progress (Outline)
@@ 1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2=0$, according to Galois theory of 1830, 1832.
Nevertheless, in his constructive proof of the Fundamental Theorem of Algebra in [120], Herman Weyl has rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough.
This solved the following Problem 0 of modern polynomial root-finding:
@@
a positive error tolerance $\epsilon=1/2^{b}>0$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle.
OUTPUT: approximations within $\epsilon$ to every zero of $p(x)$ (aka every root of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$. lying in the domain $\mathbb{D}$ such that
$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm if}~{}z_{j}\in\mathbb{D}.$ (1.2)
@@
[Approximation of all $d$ roots.] For a real
$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.3)
(see [39] for simple expressions for such bound $R$) and $\mathbb{D}$ denoting the disc $D(0,R):=\{x:~{}|x|\leq R\}$, Problem 0 amounts to approximation of all $d$ roots $z_{j}$ within $\epsilon$.
@@
[Numerical real root-finding.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots [50, 29]. Now consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon$-neighborhood of a fixed real line segment. This task addresses approximation of both real and near-real roots, that is, their perturbations due to small input and rounding errors. The known real polynomial root-finders (cf. [63, Chs. 2 and 3]) ignore this important issue.
@@
Solution of Problem 0 can be extended to the eigenvalue problem for matrices and matrix polynomials (see our Secs. 1.3 and 1.5 and Part (C) of Sec. 5.4) and is involved into the solution of a multivariate polynomial system of equations [21]; both of them are much studied areas of modern computations. [65, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics. Furthermore, Problem 0 for a domain $\mathbb{D}$ containing all roots is closely linked to polynomial factorization (see Secs. 22 and 23), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [121, 11, 5, 26, 27, 118].
@@
Hundreds of efficient polynomial root-finders have been proposed [63, 65] and keep appearing (see [97, 15, 30, 54, 59, 99, 89, 16, 17, 49, 91, 92, 58, 45, 47, 64, 48, 44, 96], and the bibliography therein). In [105] Smale proposed to study Problem 0 in terms of its arithmetic complexity and computational precision. Together they would define its bit operation complexity, but Smale deduced no complexity estimates. [106] Schönhage did this in [106]. His much involved root-finder, presented with terse complexity analysis on 46 pages, approximates all $d$ roots by using $\tilde{O}(bd^{3})$ bit operations, that is, $bd^{3}$ bit operations up to a factor in $bd$ [^2] Error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable, and assume that $b\geq d$ throughout our paper; alternatively one can replace $b$ with $b+d$ in our upper bounds on the computational cost.
@@
His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [78] (also see its journal versions [79, 84] and surveys [80, 81]). Both upper bound in $\tilde{O}(bd^{2})$ and lower bound of $0.25(d+1)bd$ bit operations of [78, 84] decrease by a factor of $d$ for polynomials of Def. 22.1 of Sec. 22, whose all zeros and tiny clusters of zeros are pairwise sufficiently well isolated (see Thms. 22.1 and 22.4 and Cor. 22.4 in Sec. 22).
@@ 1.2 Reduction to the case of the unit disc
Hereafter $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at a complex point $c$ and its boundary circle, respectively, and Problem $\widehat{0}$ denotes Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
@@
(i) Two ways from the roots in the unit disc to all roots. Given a value $R$ of (1.3), we can scale the variable, $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$, and then approximate every root $x_{j}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
Here is Schönhage’s alternative reduction in [106] of Problem 0 for all $d$ roots to Problem $\widehat{0}$: Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
@@
(ii) From roots in the unit disc to roots in a bounded domain. Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.4), and finally (d) use the map (1.4) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$. In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$. In Sec. 2.1 we curtail such an extraneous work (see Prop. 2.1 and Remark 2.1).
1.3 Black box root-finders and approximation of matrix eigenvalues
A nontrivial variant of Newton’s iterations by Louis and Vempala at FOCS 2016 [59] approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.3). While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, Louis and Vempala in [59] never involve them, working for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation, and then extend their root-finder to efficient approximation of the eigenvalues of a matrix. Next we state the relevant tasks.
@@ 1.3 Black box root-finders and approximation of matrix eigenvalues
Problem 2: Approximation of the Eigenvalues of a Matrix. [^3] As in [59], the problem covers or can be extended from a matrix to a polynomial matrix, aka matrix polynomial.
INPUT: $\epsilon=1/2^{\tilde{b}}>0$, an integer $d>0$, a $d\times d$ matrix $M$, and a domain $\mathbb{D}$ on the complex plane.
OUTPUT: Approximations within $\epsilon$ to every eigenvalue of $M$ lying in the domain $\mathbb{D}$.
@@
Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of Roots and Eigenvalues. Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) that cover all roots or eigenvalues lying in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
@@ 1.4.1 Why we should use them
Namely, the algorithm of Fiduccia [32] and Moenck and Borodin [67] (see [12, Sec. 4.5] for historical account) evaluates a polynomial of degree $m$ at $m$ points in $O(m\log^{2}(m))$ ops if its coefficients are given, versus $2m^{2}$ ops involved in Horner’s algorithms. Numerically stable algorithms of [64, 44] yield such acceleration also in terms of bit operation cost.
One can perform black box evaluation at $q>d$ points $x_{1},\dots,x_{q}$ and then interpolate to $p$. The bit operation cost of interpolation, however, is unbounded unless the $q$ points are well isolated from the roots, but it is a challenge to yield such isolation. Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1, 1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots as well as where a polynomial $p$ can be evaluated fast because root-finding for a polynomial is routinely reduced to its evaluation at sufficiently many points [63, 65].
For example, one only needs $O(\log(d))$ ops for evaluation of a Mandelbrot-like polynomial $p:=M_{k}(x)$, where $M_{0}:=x,~{}M_{i+1}(x):=xM_{i}^{2}(x)+1$, $i=0,1,\dots,k$, $d=2^{k}-1$, or the sum of a bounded number of shifted monomials, such as $p:=\alpha(x-a)^{d}+\beta(x-b)^{d}+\gamma(x-c)^{d}$ for six constants $a$, $b$, $c$, $\alpha$, $\beta$, and $\gamma$. For comparison, if we compute the coefficients of such a polynomial and then use them to evaluate it at a point $x$ by applying Horner’s algorithm, then we would need $2d$ ops, and this bound is optimal (see [72, 53] or [112, Section “Pan’s method"]).
In computational practice numerical stability of polynomial evaluation is highly important, and it is ensured when a polynomial is represented in the Bernstein basis [57]. Transition to it from the monomial basis is costly but is not needed where a root-finder works for a black box polynomial.
One can benefit from black box root-finders even where an input polynomial is given with its coefficients.
(i) Application of a black box root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ matrix $M$ only involves the values of this polynomial but not its coefficients, whose overall size exceeds the size (a norm) of the matrix $M$ by a factor of $d$. Hence extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [106, Sec. 21].
@@
(iii) Deflation of a factor of a polynomial is a well-known cause of the coefficient swell, e.g., for degree 300 factors of $x^{1000}+1$, which blows up precision and bit complexity of computations. For an exception, one can readily deflate a small degree factor $f(x)$ and approximate its zeros, but the coefficients of the high degree factor $g(x)=p(x)/f(x)$ still tend to blow up. We, however, avoid their computation when we apply a black box root-finder to the polynomial $g(x)$ represented by the pair of $p(x)$ and $f(x)$.
1.4.2 Progress at FOCS 2016 and its limitations